OpenStudy (anonymous):

5. Prove that if x an irrational number and y is rational number, then x+y is an irrational number.

OpenStudy (anonymous):

Let x+y be rational, then \[x+y=\frac{ p }{q },(say),wher p and q are integers.\] \[x=\frac{ p }{ q }-y=\frac{ p-qy }{y }=\frac{ s }{ t },(say)where s and t are integers.\] or x is a rational number. which is a contradiction. because x is a rational number. (given) Hence our assumption is wrong. or x+y is an irrational number.

OpenStudy (anonymous):

Sum of a rational and an irrational is always irrational. Proof: Let the rational number be of the form pq, where p∈ℤ and q∈ℤ∖{0} while the irrational number be r. If r+pq is a rational, then we have that r+pq=ab for some a∈ℤ and b∈ℤ∖{0}. This means that r=ab−pq=aq−bpbq where aq−bp∈ℤ and bq∈ℤ∖{0}. This contradicts the fact that r is irrational. Hence, our assumption that r+pq is a rational is false. Hence, r+pq is a irrational.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

give me a medal :)